OpenStudy (anonymous):

Lisa is now six years older than cheryl, in 2 years lisa will be twice as old as cheryl how old is lise now? Got 2x+8=x+2?

OpenStudy (anonymous):

present age of Cheryl is x and that of Lisa's present age is x+6 after two yrs age of Cheryl is x+2 and that of Lisa is x+6+2 =x+8 Lisa being elder than Cheryl and twice of age of Cheryla after 2 yrs so we can write x+8 =2(x+2)

OpenStudy (anonymous):

so now you may find x